OpenStudy (anonymous):

Please help! Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. C = 67°, a = 21, c = 20

OpenStudy (raden):

The Law of Sines : a/sinA = c/sinC

OpenStudy (raden):

or sinA = (a sinC / c)

OpenStudy (raden):

just put the values of them, giving us : sinA = 21*sin67° / 20 = .... (use the calculator to get sin67°)

OpenStudy (anonymous):

I got .966

OpenStudy (raden):

finally, after u got the value of sinA, take the arcus sin to get the measure of A

OpenStudy (anonymous):

do you mean arcos?

OpenStudy (raden):

sinA = 0.966 A = arc(sin(0.966) = ... (use the calculator again :))

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

umm, do I use arcos and then sin?

OpenStudy (raden):

it is the value of sines, why we use arcos ??

OpenStudy (anonymous):

I got 75 using arcsin(.966) but i dont think thats right

OpenStudy (raden):

yes, that's right (nearest)

OpenStudy (raden):

but, dont ferget the identity of : sin (180 - x) = sin x so, sin75 = sin (180-75) = sin105 therefore there are 2 values of A satisfies, are A = 75 or A = 105 degrees

OpenStudy (raden):

* forget

OpenStudy (anonymous):

so how do I decide which answer is correct? A = 75.1°, B = 37.9°, b = 13.3; A = 104.9°, B = 8.1°, b = 3.1 A = 75.1°, B = 37.9°, b = 30; A = 104.9°, B = 8.1°, b = 30

OpenStudy (raden):

ah, ur option wants the nearest of A=75.1

OpenStudy (raden):

so, if there are 2 measures of A=76.1 and C=67, to get B : B=180-(76.1+67) = 180 - 143.1 = 36.9 degrees

OpenStudy (raden):

looks no option there, haha

OpenStudy (raden):

oppss, i do mistake it should 75.1 not 76.1 :)

OpenStudy (raden):

let's try again : B = 180 - (75.1+67) = 180 - 142.1 = 37.9 degrees

OpenStudy (raden):

and because there are 2 values of A, are A=75.1 and from sin (180 - x) = sin x so, sin75.1 = sin (180-75.1) = sin104.9 thus, A=104.9 deg

OpenStudy (raden):

to get B : B=180-(67+104.9) B=180-171.9 = 8.1 degrees

OpenStudy (raden):

SOLVED

OpenStudy (anonymous):

but does b= 13.3 or 30?

OpenStudy (raden):

ah, the length of b right ? sorry, it's left

OpenStudy (raden):

use the law of sines again a/sinA = b/sinB or b = a*sinB/sinA = ... just put the values of a, sinB, and sinA (dont forget there are are 2 the values of A, so will get 2 values of B too)

OpenStudy (raden):

try it ...

OpenStudy (anonymous):

whats the value of a again?

OpenStudy (raden):

it's given in ur question above :P

OpenStudy (anonymous):

oh 21 okay

OpenStudy (raden):

a is the length of side, but A is the measure of angle A they are different

OpenStudy (anonymous):

I understand, I got b = 13.3

OpenStudy (anonymous):

So it would be the first option i had given

OpenStudy (raden):

the 2nd ?

OpenStudy (anonymous):

no, b = 30 in the 2nd option...

OpenStudy (raden):

nevermind, exactly the 2nd is 3.1 right ?

OpenStudy (raden):

i mean the value for b

OpenStudy (anonymous):

oh yes, thats correct

OpenStudy (raden):

okay, game is over

OpenStudy (anonymous):

can you help me answer a much simpler question quickly/

OpenStudy (raden):

the new question ?

OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. A = 61°, a = 23, b = 24